OpenStudy (anonymous):

if f(x)=sinx and g(x)=4x-3, find f(g(x))

OpenStudy (anonymous):

To find f(g(x)) we have to go to the f(x) and plug g(x) instead of x. f(g(x))=sin(g(x))=??

OpenStudy (anonymous):

so you get sin (4x-3) ....im just so confused after that. im not good with trig

OpenStudy (lgbasallote):

isnt that the final answer already? o.O

OpenStudy (anonymous):

oh, is it? hahah i wasn't sure

OpenStudy (anonymous):

@rjeffries96 that's your final answer

OpenStudy (lgbasallote):

or do you want to express it in something like \[\sin (a + b) = \sin a \cos b + \sin b \cos a\] i prefer writiing it as is

OpenStudy (anonymous):

yeah i thought about that also...i have no idea which to do. my calc teacher is giving us this question online so i dont know how far to take it

OpenStudy (lgbasallote):

calculus huh

OpenStudy (lgbasallote):

the bring it farther than algebra haha

OpenStudy (anonymous):

ok haha..thank you, i really appreciate it :)

OpenStudy (anonymous):

@rjeffries96 , the simplest solution is usually the acceptable solution, so I think you should stick to your first answer

OpenStudy (lgbasallote):

just doing my avatar duty

OpenStudy (anonymous):

maybe i could just type in f(g(x))=sin(4x-3)=sin4xcos3+sin3cos4x?